John has dimes and quarters worth $6.50 he has two less quarters than dimes. how many of each type does he have?

Answer by checkley79(3341)

.10D+.25Q=6.50
D=Q+2
.10(Q+2)+.25Q=6.50
.10Q+.20+.25Q=6.50
.35Q=6.50-.20
.35Q=6.30
Q=6.30/.35
Q=18 QUARTERS.
D=18+2
D=20 DIMES.
PROOF:
.10*20+.25*18=6.50
2.00+4.50=6.50
6.50=6.50